OpenStudy (anonymous):

Help please i will medal and fan: Find the distance between the points on the coordinate plane. Points B and (–1, 2) A. 3 B. 4 C. 5 D. 6

OpenStudy (anonymous):

Heres the graph

hartnn (hartnn):

can you tell me co-ordinate of point B first ?

OpenStudy (anonymous):

(5 2)

hartnn (hartnn):

thats correct, now you have both points, use the distance formula!

hartnn (hartnn):

Distance between points (x1,y1) and (x2,y2) is \(\huge d=\sqrt{(x_1-x_2)^2+(y_1-y_2)^2}\) does this help ??

OpenStudy (anonymous):

Um?

OpenStudy (anonymous):

Not really :/ But thanks

hartnn (hartnn):

the 2 points are x1, y1 = -1,2 x2, y2 = 5,2 so just plug in x1 = -1 y1 = 2 x2 = 5 y2=2 in that formula!

OpenStudy (anonymous):

ok.....

OpenStudy (anonymous):

..

OpenStudy (anonymous):

I have no idea what u said but ill medal and fan

hartnn (hartnn):

\(\huge d=\sqrt{(x_1-x_2)^2+(y_1-y_2)^2} \\ \huge d=\sqrt{(-1-5)^2+(2-2)^2}\)

hartnn (hartnn):

what is -1-5 = ... ?

OpenStudy (anonymous):

4

hartnn (hartnn):

not actually -1 -5 = - (1+5) = -6 got this ?

OpenStudy (anonymous):

Yes Thank you very mcuh

hartnn (hartnn):

\( \huge d=\sqrt{(-1-5)^2+(2-2)^2} \\ \Large d = \sqrt{(-6)^2+0} = \sqrt{36}=6\) thats how we solve distance formula related problems :)

OpenStudy (anonymous):

Ok ill write that on my notebook :) Thank you.

hartnn (hartnn):

welcome ^_^